OpenStudy (anonymous):

find the four arithmetic means between -21 and -36

OpenStudy (solomonzelman):

there are 4 arithmetic means between -21 and -36, so we can say that \(\large\color{blue}{ a_{n}=-21 }\) and \(\large\color{blue}{ a_{n+5}=-36 }\)

OpenStudy (solomonzelman):

So, \(\large\color{blue}{ a_{n+5}-a_{n~}=5d }\)

OpenStudy (solomonzelman):

-36-(-21)=5d → 5d =-15 → d=-3

OpenStudy (solomonzelman):

So if d=-3 and the first term is -21, can you will in the missing means ?